# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
$=\int_{[0,\varepsilon_{0}]\times B_{N}(0)}\bigg[v^{\infty}_{t}(s,y)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(y)D^{2}_{x}v^{\infty}(s,y)\right)+b^{\varpi}(y)D_{x}v^{\infty}(s,y)+r^{\varpi}(s,y)\bigg]p^{x}(s,y)dyds$
$\leq\int_{[0,\varepsilon_{0}]\times B_{N}(0)}\bigg[v^{\infty}_{t}(s,y)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(y)D^{2}_{x}v^{\infty}(s,y)\right)$
$\qquad\qquad\qquad\quad+b^{\pi^{\infty}}(y)D_{x}v^{\infty}(s,y)+r^{\pi^{\infty}}(s,y)+2Cs^{\alpha/2}\bigg]p^{x}(s,y)dyds$
$=\lim_{k\to\infty}\int_{[0,\varepsilon_{0}]\times B_{N}(0)}\bigg[v^{n_{k}}_{t}(s,y)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(y)D^{2}_{x}v^{n_{k}}(s,y)\right)$
$\qquad\qquad\qquad\qquad\quad+b^{\pi^{n_{k}}}(y)D_{x}v^{n_{k}}(s,y)+r^{\pi^{n_{k}}}(s,y)+\lambda_{n_{k}}\delta(s)\mathcal{H}(\pi^{n_{k}}(x))+2Cs^{\alpha/2}\bigg]p^{x}(s,y)dyds$
$\leq\lim_{k\to\infty}\int_{[0,\varepsilon_{0}]\times B_{N}(0)}\bigg[v^{n_{k}}_{t}(0,y)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(y)D^{2}_{x}v^{n_{k}}(0,y)\right)$
$\qquad\qquad\qquad\qquad\quad+b^{\pi^{n_{k}}}(y)D_{x}v^{n_{k}}(0,y)+r^{\pi^{n_{k}}}(0,y)+\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}}(x))+3Cs^{\alpha/2}\bigg]p^{x}(s,y)dyds$
$=\int_{[0,\varepsilon_{0}]}3Cs^{\alpha/2}p^{x}(s,y)dyds=\frac{3C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2},$
@@
where the first inequality follows from (4.13), the third equality follows from (4.7), Lemma 4.1 part (i) and the fact that $\lim_{k\to\infty}\|\lambda_{n_{k}}\mathcal{H}(\pi^{n_{k}}(x))\|_{L^{\infty}(\mathbb{R}^{d})}=0$, the second inequality follows from (4.14), and the last equality follows from (4.2) with $t=0$. Thanks to (4.15) and (4.16), it holds that
@@
$J^{\varpi\otimes_{\varepsilon_{0}}\pi^{\infty}}(x)-J^{\pi^{\infty}}(x)=\mathbb{E}_{x}\left[v^{\infty}({\varepsilon_{0}},X^{\varpi}_{{\varepsilon_{0}}})+\int_{0}^{\varepsilon_{0}}r^{\varpi}(s,X^{\varpi}_{s})ds\right]-J^{\pi^{\infty}}(x)\leq\frac{3C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2}+\epsilon.$
By the arbitrariness of $\epsilon$, we have $J^{\varpi\otimes_{\varepsilon_{0}}\pi^{\infty}}(x)-J^{\pi^{\infty}}(x)\leq\frac{3C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2},$ and hence $\pi^{\infty}$ fulfils the definition of equilibrium in (2.4). ∎
From the previous proof, we reveal a “weak type” equilibrum HJB system as a characterization for the value function and its associated equilibrium. In particular, we have the following weaker sufficient condition for the existence of equilibrium without requiring the conventional strong regularity assumptions.
Corollary 4.1.
Suppose $\pi^{*}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ is an admissible relaxed control. If $u\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ and $\sup_{x\in\mathbb{R}^{d}}\|u(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d}}<\infty$, and there exists $\varepsilon_{0}$ such that $u$ satisfies the following PDE system in distribution:
$Ct^{\alpha/2}\geq$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D^{2}_{x}u(t,x)\right)$
$+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(t,x)+r(t,x,a)\right]\varpi(da)\right\},\;\text{a.e on $[0,\varepsilon_{0}]\times\mathbb{R}^{d}$},$ (4.17)
$0=$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}u(t,x)\right)+b^{\pi^{*}}(x)D_{x}u(t,x)+r^{\pi^{*}}(t,x)\quad\text{a.e. on $\mathbb{T}\times\mathbb{R}^{d}$},$ (4.18)
then $u=V^{\pi^{*}}$ on $\mathbb{T}\times\mathbb{R}^{d}$ and $\pi^{*}$ is an equilibrium.

### 04 Apr 2026 v2

@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
$\leq\int_{0}^{\varepsilon_{0}\wedge\rho_{N}}\bigg[\partial_{t}V^{\pi*}(s,X^{\varpi}_{s})+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\varpi}_{s})D^{2}_{x}V^{\pi*}(s,X^{\varpi}_{s})\right)$
$\qquad\qquad\qquad\quad+b^{\pi^{*}}(y)D_{x}V^{\pi*}(s,X^{\varpi}_{s})+r^{\pi^{*}}(s,X^{\varpi}_{s})+2Cs^{\alpha/2}\bigg]ds$
$\leq\int_{0}^{\varepsilon_{0}}2Cs^{\alpha/2}ds=\frac{2C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2},$
@@
where the first inequality follows from (4.19), and the second last equality follows from that $V^{\pi^{*}}$ satisfies (4.2). Thanks to (4.21) and (4.22), it holds that
@@
$J^{\varpi\otimes_{\varepsilon_{0}}\pi^{*}}(x)-J^{\pi^{*}}(x)=\mathbb{E}_{x}\left[V^{\pi*}({\varepsilon_{0}},X^{\varpi}_{{\varepsilon_{0}}})+\int_{0}^{\varepsilon_{0}}r^{\varpi}(s,X^{\varpi}_{s})ds\right]-J^{\pi^{*}}(x)\leq\frac{2C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2}+\epsilon.$
By the arbitrariness of $\epsilon$, we have
$J^{\varpi\otimes_{\varepsilon_{0}}\pi^{*}}(x)-J^{\pi^{*}}(x)\leq\frac{2C}{1+\alpha/2}\varepsilon_{0}^{1+\alpha/2},$
and hence $\pi^{*}$ fulfills Definition 2.1.